OpenStudy (anonymous):

Alright.. So I Thought I Knew How To Do This. Turns Out I Only Understand That Certain Problem.. Just Not THIS One: 5 ÷ 10/3

hero (hero):

\[5 \dot\ \frac{3}{10} = \frac{15}{10} = \frac{3}{2}\]

OpenStudy (anonymous):

Thanks.